You are asked to design a system to degrade phenyl acetate (CH3COOC6H5). With regard to the impact of hydrolysis on this compound, would you recommend a neutral (pH=7), acidic (pH=4) or basic (pH=10) environment? Calculate the overall hydrolysis rate for each of these three conditions. (10 points)
Answer #1

The hydrolysis of esters is catalyzed by either an acid or a base.so under ph =4 and ph=10 hydrolysis will take place.

CH3COOC6H5 + H3O+ =CH3COOH + C6H5OH

SINCE water is taken in excess, rate will depend on ester cinc only, so rate =k(CH3COOC6H5)

Under basic conditions

CH3COOC6H5 + NaOH =CH3COONA+ C6H5OH

RATE =K1(Ch3COOC6H5) (OH-)

Under neutral ph hydrolysis will not take place.

I hope it helps you, please give me a positive feedback ?